OpenStudy (anonymous):

can some one help me find the y-intercept of the graph of this equation: 4y + 8x = 24

OpenStudy (anonymous):

y=mx+b so b is the y-intercept

OpenStudy (anonymous):

m=−2,b=6 where m is the slope and b is the y-intercept

OpenStudy (anonymous):

4y + 8x = 24 = 4y = -8x + 24 can you finish

OpenStudy (anonymous):

Let's solve for x. 4y+8x=24 Step 1: Add -4y to both sides. 8x+4y+−4y=24+−4y 8x=−4y+24 Step 2: Divide both sides by 8. 8x8=−4y+248 x=−12y+3 Answer: x=−12y+3